The future Teachers of America Club at Paint Branch High School is maaking a healthy trail mix to to sell during lunch. The mix will have three times the number of pounds of raisins as sunflower seeeds. Sunflower seeds cost $4.00 per pound, raisins cost $1.50 per pound. If the group has $34.00 to spend on the raisins and sunflower seeds , how many pounds of each should they buy.
(please solve this problem by using substution)

Answer by RAY100(1637)

Let R= # lbs of raisins,,,,,@ $1.50 per #
Let S= # lbs of sunflower seeds,,,,@$4.00 per #
given,,,3R=S
given,,,, 4S+1.5R=34
4(3R) +1.5R=34
12R+1.5R=34
13.5R=34
R= 2.52 #
but,,,,3R=S=3(2.52)=7.5#
check
4(7.5)+1.5(2.52)=34
30.22+3.78=34
34=34,,,,ok